# CAN-STRESS: A Real-World Multimodal Dataset for Understanding Cannabis Use, Stress, and Physiological Responses

### v1

@@ VI Conclusion
In this paper, we introduced CAN-STRESS, a multimodal dataset designed to advance research into the physiological and behavioral effects of cannabis use in real-world settings [^1] https://zenodo.org/records/14842061. By integrating self-reported data on key daily activities, such as cannabis use, exercise, sleep patterns, and subjective stress ratings, with high-resolution physiological signals collected via a wearable wristband, CAN-STRESS provides a unique opportunity to examine the relationship between subjective experiences and objective measurements. Addressing the limitations of lab-based studies, CAN-STRESS offers an ecologically valid foundation for exploring stress regulation, activity recognition, and other health-related research domains. We encourage researchers to utilize CAN-STRESS to drive advancements in behavioral science, wearable computing, and medical diagnostics, paving the way for new insights into cannabis use and its broader implications.

### v2

@@ IV Conclusion
In this paper, we introduced CAN-STRESS, a multimodal dataset designed to advance research into the physiological and behavioral effects of cannabis use in real-world settings [^1] https://zenodo.org/records/14842061. By integrating self-reported data on key daily activities with high-resolution physiological signals collected via a wearable wristband, CAN-STRESS provides a unique opportunity to examine the relationship between subjective experiences and objective measurements. Addressing the limitations of lab-based studies, CAN-STRESS offers an ecologically valid foundation for exploring stress regulation, activity recognition, and other health-related research domains. We encourage researchers to utilize CAN-STRESS to drive advancements in behavioral science, wearable computing, and medical diagnostics.